OpenStudy (anonymous):

The line of symmetry of the parabola whose equation is y = ax2 - 4x + 3 is x = -2. What is the value of "a"? Step by step please

sam (.sam.):

You know completing the square?

OpenStudy (anonymous):

Kind of

sam (.sam.):

\[ax^2 - 4x + 3\] \[a(x^2-\frac{4x}{a})+3\] \[a(x-\frac{2}{a})^2-(\frac{4x^2}{a^2})(a)+3\] ------------------------------------------------ \[x=\frac{2}{a}\] a should be -1